yourbxbygirl:

i need help with a few math questions- Which of the following fractions is equal to a terminating decimal? 2/7 5/8 1/3 2/9

men:

A terminating decimal, true to its name, is a decimal that has an end. so there fore where figuring out which one has a end to it i think. so between A B C and D which one would the answer be we can try to use process of elimination. so basically i think you have to wright our the full decimal on each one and which ever one has a end is the answer. thats all ive got.

men:

out*

men:

so we right out the decimals 2/7 = 0.28571428571 5/8 = 0.625 1/3 = 0.3333 2/9 = 0.22222222 i think thats them in decimal form so thats the first step next you use elimination and find your answer